2 cars are 500feet apart, traveling toward each other at 40mph.How much time does it take to be even with each other?
Answer by mananth(16949)

Towards each other so add up speeds.
80 mph
500 ft = 500/5280 miles
t= d/r
t=(500.5280) /(80)
500 / (5 280 * 80) = 0.00118371212
0.00118371212* 3600 = 4.26 seconds
( 1 hour = 3600 seconds)